# BRcal: An R Package to Boldness-Recalibrate Probability Predictions

### v2

@@ 3.6.2 Color Contour Plot
The second visualization type is a color contour plot showing how the posterior model probability, $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$, changes as a function of $\delta$ and $\gamma$. This visualization is implemented in the plot_params function and utilizes base R graphics and the fields package [28]. The resulting plot shows the posterior model probability surface over a $k$ by $k$ grid of $\delta$ and $\gamma$ values using a color scale to differentiate high and low values of $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$. Additionally, users have the option to add contours at specified levels of calibration using t_levels. The goal of this is to visualize different values of $t$ at which a user may want to boldness-recalibrate, but does not perform boldness-recalibration nor identify $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$.
@@
Some care is required for the selection of the range of $\delta$ and $\gamma$ via arguments dlim and glim, and the size of $k$ via argument k, which together determine the fineness of your grid of parameter values. Small values of k or misspecified parameter bounds may cause difficulties in detecting the region of non-zero $P(M_{c}|\mathbf{y},\mathbf{x}^{\prime}_{ij})$s. Large values of k may cause a computational bottleneck as computation time increases with k. Appendix B presents recommendations for specifying k, dlim, and glim. Once users have settled on values of these arguments, you can specify return_z=TRUE to return the underlying matrix of posterior model probabilities to avoid recalculation of these values. This matrix can then be reused by passing it to argument z, indicating to plot_params() that no recalculations are needed.
@@
4 Application
In this section we demonstrate BRcal using a case study related to the foreclosure risk modelling presented by Keefe et al. [30]. Note that in all places in the output where ... appears, this indicates that the output was truncated for readability.
4.1 Foreclosure Monitoring Predictions Case Study
This example uses the foreclosure dataframe built into the BRcal package. We begin by loading the BRcal package and foreclosure data and printing the first few rows of the data via the code below.
@@ 4.1 Foreclosure Monitoring Predictions Case Study
4.1.1 Assessing Calibration
With the data loaded, we can assess the calibration of the foreclosure predictions. First we will use the bayes_ms function to use the approximate Bayesian model selection approach.
@@ 4.1.1 Assessing Calibration
Notice from the output that bayes_ms() returns seven quantities in the form of a list. The prior model probability for $M_{c}$ is returned in Pmc to remind users what value was specified (0.5 by default). List items BIC_Mc and BIC_Mu in (6) and (7) are the BICs for the calibrated model and uncalibrated model, respectively. The Bayes factor comparing $M_{u}$ to $M_{c}$ in (5) is returned in BF and posterior_model_prob is the posterior model probability of calibration from (3). The MLEs for the vector x are returned in MLEs. Additionally, the list returned by the call to optim() to get the MLEs is returned in sublist optim_details.
@@
The results of bayes_ms() and llo_lrt() both indicate that the foreclosure predictions from the model trained in 2005 are poorly calibrated, as the posterior model probability and the p-value are near 0.
4.1.2 MLE Recalibration
Next, we demonstrate how to use mle_recal() to recalibrate the foreclosure predictions such that calibration is maximized. By default, mle_recal() returns a list with (i) the vector of maximally calibrated predictions in probs, (ii) the MLEs in MLEs and (iii) the details from optim() in optim_details. Should a user only want the vector of MLE-recalibrated probabilities, they can set the option probs_only = TRUE.
> mle <- mle_recal(foreclosure\$x, foreclosure\$y, probs_only = FALSE)
> mle
\$probs
[1] 0.4557362 0.8600894 0.4905113 0.8812196 0.7491834 0.9415870
[7] 0.9247097 0.5139758 0.9206578 0.6645239 0.5170544 0.8372189
[997] 0.8972962 0.8860154 0.3612379 0.9433734
[ reached getOption("max.print") -- omitted 4000 entries ]
\$MLEs
[1] 11.109412  1.270519
\$optim_details
\$optim_details\$par
[1] 11.109412  1.270519
\$optim_details\$value
[1] 2071.375
\$optim_details\$counts
function gradient
      85       NA
\$optim_details\$convergence
[1] 0
\$optim_details\$message
NULL
4.1.3 Boldness-Recalibration
Now we boldness-recalibrate using the brcal() function. By default, brcal() performs 95% boldness-recalibration and prints each step of the optimization process from nloptr(). To suppress all output from nloptr(), users can set print_level=0. Other options to minimize output can be found in the nloptr() documentation and can be passed to brcal() via print_level. Below we show a example where the full output from the first and last steps are show, the in between stages are truncated for readability.
> br95 <- brcal(hockey\$x, hockey\$y)
@@ 4.1.3 Boldness-Recalibration
Notice the large amount of output from nloptr() which is returned by brcal() in nloptr. As with any optimization routine, we include this copious output to allow users to check that the routine converged properly. While this output shows the value achieved optimal values of the parameters and objective function, we choose to return these separately in the list items BR_params and sb, respectively for easier extraction. The resulting set of boldness-recalibrated probabilities are also returned for easier extraction.
We enforce different restrictions on calibration by changing the value of parameter t. The two lines of code below perform 90% and 80% boldness-recalibration. We also demonstrate suppressing the output from nloptr by setting print_level=0.
br90 <- brcal(foreclosure\$x, foreclosure\$y, t=0.9, print_level=0)
br80 <- brcal(foreclosure\$x, foreclosure\$y, t=0.8, print_level=0)
Note that by leveraging the gradient information and efficient nloptr package as described in Section 3.3, each of these calls to brcal() for the foreclosure data takes approximately 60 seconds or less to run.
4.1.4 LLO-Adjustment
The LLO() function allows users to directly LLO-adjust a vector of probability predictions, passed via x, by shift parameter delta and scale parameter gamma. While most users of this package may not use the LLO() function directly, there are some cases where users may want to do so. One example of this could be when you have MLEs for $\delta$ and $\gamma$ based on historical data and want to LLO-adjust unlabeled out of sample predictions. A user could use the LLO function to LLO-adjust the new set of predictions by passing them to function argument x, and the MLEs for $\delta$ and $\gamma$. Another use case could be when you have boldness-recalibration parameter estimates from brcal() but did not save the boldness-recalibrated probabilities. Users could LLO-adjust their probabilities using the boldness-recalibration parameter estimates using LLO()without needing to re-compute them using brcal(), thus avoiding unnessesary computation time. This is demonstrated below with the 95% boldness-recalibration estimates saved in section 4.1.3.
> LLO(foreclosure\$x, delta=br95\$BR_params[1], gamma=br95\$BR_params[2])
   [1] 0.4069886 0.8627809 0.4448838 0.8856910 0.7381995 0.9483221
   [7] 0.9313465 0.4707687 0.9271990 0.6411276 0.4741822 0.8375914
 [997] 0.9028353 0.8908336 0.3073575 0.9500874
 [ reached getOption("max.print") -- omitted 4000 entries ]
Notice in the output above that we get the same boldness-recalibrated probabilities as we did using brcal() in section 4.1.3.
4.1.5 Visualizations
First we will use lineplot to visualize how MLE-recalibration, t=95, 90, and 80% boldness-recalibration changed the individual probabilities themselves. To do so, we pass the vector of the original probability predictions and the corresponding outcomes along with the levels of boldness-recalibration we want to visualize using t_levels. The code below generates the lineplot that is shown on the left in Figure 2. Given the size of this data, the plot may take a few minutes to generate. In practice, users can leverage the strategies discussed in section 3.6.1 for reducing computation time.
> lp <- lineplot(foreclosure\$x, foreclosure\$y, t_levels=c(0.95, 0.9, 0.8),
                 return_df=TRUE)
> lp\$plot
While it can be useful to visualize all of these sets at the same time, users may want to focus on how emboldening affects predictions after they have already been MLE-recalibrated. The following line of code uses the dataframe returned from the first call to lineplot() to plot only the MLE-recalibrated and 95% boldness-recalibrated sets as see in the right plot in Figure 2.
> lineplot(df=lp\$df, plot_original=FALSE, t_levels=c(0.95))
Figure 2: Demonstration of the lineplot() function visualizing how predictions change from varying recalibration strategies. On the left plot, the first column of points is the original set of probability predictions. The second column are the MLE-recalibrated predictions (i.e., after recalibrating with $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$). The last three columns are the predictions after 95%, 90%, and 80% boldness-recalibration (i.e. $\mathbf{x}_{0.95}$, $\mathbf{x}_{0.90}$, and $\mathbf{x}_{0.80}$) respectively. On the right plot, the columns are the MLE-recalibrated and 95% boldness-recalibrated probabilities. A line is used to connect each original prediction to where it ends up after each recalibration procedure. Blue points and lines correspond to predictions for which there was a foreclosure (i.e. “events”). Red corresponds to transactions in which there was no foreclosure (i.e. “non-events”). Achieved $P(M_{c}|\mathbf{y})$s are reported on the x-axis in parentheses.
There are a few key takeaways from the visualizations in Figure 2. First, the original foreclosure predictions are relatively accurate in terms of classification accuracy as more of the observations are colored blue (corresponding to foreclosure) for higher probabilities and more are colored red (corresponding to non-foreclosure) for lower probabilities. In fact, these predictions achieve an area under the ROC curve (AUC) of 0.8. Second, the original predictions are substantially biased, as they consistently underestimating the rate of foreclosure and MLE-recalibration corrects this bias. This is mostly likely due to using a model trained on pre-recession data from 2005 to estimate foreclosure rates for transactions from post-recession in 2010. Finally, these predictions can also benefit from being spread out under boldness-recalibration, as their spread has increased from the original set to the 95% boldness-recalibrated set. This is the maximum boldness achievable while maintaining a calibration probability of 0.95, subject to classification accuracy.
Next, we will use the plot_params() function to see the posterior model probability surface. As described in 3.6, we need to specify bounds for $\delta$ and $\gamma$ over which we want to plot and the density of the grid of values. Arguments dlim and glim take the bounds for $\delta$ and $\gamma$ over which the user wishes to plot. Argument k is used to specify a value for k. In the code below, we examine k=200 values of $\delta\in[8.5,15]$ and 200 values of $\gamma\in[1.05,1.5]$. Since we are interested in 95, 90, and 80% boldness-recalibration, we will also use t_levels=c(0.95, 0.9, 0.8) to specify that we want contours drawn at those three levels. Additionally, since this plot uses base R graphics, we can add points denoting the boldness-recalibration parameters and MLEs using points(). Recall, these parameter estimates are saved in br95, br90, br80, and bt. The result of the code below to generate the lineplot is shown in Figure 3
@@ 4.1.5 Visualizations
Figure 3: Color contour plot demonstrating the plot_params() function. The points in white along the contours denote the boldness-recalibration parameters. The $\times$ denotes the MLEs.
The key takeaways from this visualization are the locations of high posterior model probability and the contours denoting the regions of interest for boldness-recalibration. This plot can aid users is validating the MLEs and boldness-recalibration estimates for $\delta$ and $\gamma$. For example, if the optimizer provides boldness-recalibration estimates that lie outside of the corresponding contour on the plot, users may need to make changes to the optimization routine to achieve convergence. Similarly, users should be concerned if the MLEs do no fall roughly in the middle of the dark red (high probability of calibration) region. In this cast study, the boldness-recalibration parameters are located along their corresponding contour. While the optimization is not restricted to searching only directly on the contour of $P(M_{c}|\mathbf{y})=t$, the optimal boldness-recalibration parameters all fall on their corresponding contours. A reason for this is that $\gamma$ is the scale parameter on the log-odds scale. Thus values of $\gamma$ that are further from 0 produce more bold predictions. Additionally, this is a direct result of the tension between calibration and boldness. As we increase $\gamma$ beyond $\hat{\gamma_{MLE}}$, i.e. more away from maximal calibration marked by the $\times$ (the location of the MLEs), we gain boldness. Naturally, this places the solution to our optimization at the boundary of the constraint which is visualized by the contours.
@@ 5 Conclusion
The BRcal package offers a variety of tools for assessing and improving probability predictions. One key contribution of this R package is the brcal() function which allows users to maximize the boldness (spread) in predictions while maintaining a specified level of calibration as measured by the posterior model probability of calibration. The other main contributions of the BRcal package are (i) MLE-recalibration, (iii) calibration assessment, and (iv) supporting visualizations. The BRcal package, with documentation and a user-friendly tutorial are available to users on CRAN and Github. Overall, the BRcal package is a useful tool to achieve more actionable predictions better decision making.
@@ Appendix A Derivation of Gradients
This appendix shows the derivation of the gradients shown in Equations (13) and (15) in Sections A.2 and A.3 using related intermediate results presented in Section A.1.

### v3

@@ 4.6.2 Color Contour Plot
The second visualization type is a color contour plot showing how the posterior model probability, $P_{\mathbf{x}^{\prime}_{ij}}(M_{c}|\mathbf{y})$, changes with $\delta$ and $\gamma$. This visualization is implemented in the plot_params() function and utilizes base R graphics and the fields package [35]. The resulting plot shows the posterior model probability surface over a $k$ by $k$ grid of $\delta$ and $\gamma$ values using a color scale to differentiate high and low values of $P_{\mathbf{x}^{\prime}_{ij}}(M_{c}|\mathbf{y})$. Additionally, users have the option to add contours at specified levels of calibration using t_levels. The goal of this is to visualize different values of $t$ at which a user may want to boldness-recalibrate, but plot_params() does not perform boldness-recalibration nor identify $\hat{\delta}_{t}$ and $\hat{\gamma}_{t}$. We also recommend this plot be used to determine if the posterior model probability surface appears to be unimodal and take on one connected region, which can be important for optimization purposes.
Arguments dlim and glim take the bounds for $\delta$ and $\gamma$ over which the user wishes to plot. Argument k is used to specify the density of the k by k grid of values. In the code below, we examine k=200 values of $\delta\in[0.5,1.5]$ and 200 values of $\gamma\in[0.25,2.75]$. Since we are interested in 95%, 90%, and 80% boldness-recalibration, we will also use t_levels=c(0.95, 0.9, 0.8) to specify that we want contours drawn at those three levels. Additionally, since this plot uses base R graphics, we can add points denoting the boldness-recalibration parameters and MLEs using points(). Recall, these parameter estimates are saved in br95, br90, br80, and bt. The code below generates the example plot in Figure 3.
> plot_params(hockey\$x, hockey\$y, dlim = c(0.5, 1.5), glim = c(0.25, 2.75),
+            k=200, t_levels = c(0.95))
> points(bt\$MLEs[1], bt\$MLEs[2], pch=4, col="white")
Figure 3: Demonstration of the plot_params() function on the hockey data.
This plot can aid users is validating the MLEs and boldness-recalibration estimates for $\delta$ and $\gamma$. For example, users should be concerned if the MLEs do no fall roughly in the middle of the dark red (high probability of calibration) region. Similarly, if the optimizer provides boldness-recalibration estimates that lie outside of the corresponding contour on the plot, users may need to make changes to the optimization routine to achieve convergence. While the optimization is not restricted to searching only directly on the contour of $P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})=t$, the optimal boldness-recalibration estimates typically fall on the upper edge of their corresponding contours. A reason for this is that $\gamma$ is the scale parameter (on the log odds scale). Thus larger values of $\gamma$ tend to produce more bold (i.e., more spread out) predictions. Additionally, this is a direct result of the tension between calibration and boldness. As we increase $\gamma$ beyond $\hat{\gamma}_{MLE}$, i.e., more away from maximal calibration marked by the $\times$ (the location of the MLEs), we gain boldness. Naturally, this places the solution to our optimization at the boundary of the constraint which is visualized by the contours. Further intuition behind this tendency is provided in Appendix F.
@@
Some care is required for the selection of the range of $\delta$ and $\gamma$ via arguments dlim and glim, and the size of $k$ via argument k, which together determine the fineness of your grid of parameter values. Small values of k or misspecified parameter bounds may cause difficulties in detecting the region of non-zero $P_{\mathbf{x}^{\prime}_{ij}}(M_{c}|\mathbf{y})$. Large values of k may cause a computational bottleneck as computation time increases with k. Appendix B presents recommendations for specifying k, dlim, and glim. Once users have settled on values of these arguments, you can specify return_z=TRUE to return the underlying matrix of posterior model probabilities to avoid recalculation of these values. This matrix can then be reused by passing it to argument z, indicating to plot_params() that no recalculations are needed. Essentially, plot_params() is using a streamlined implementation of the grid-search based approach for boldness-recalibration. The improvement made to boost speed is discussed in Appendix C.
@@
5 Foreclosure Monitoring Case Study
In this section we demonstrate BRcal using a case study related to the foreclosure risk modeling presented by Keefe et al. [38]. This example uses the foreclosure dataframe built into the BRcal package, as described in Section 4.1. We begin by loading the BRcal package and foreclosure data and printing the first few rows of the data via the code below.
@@ 5 Foreclosure Monitoring Case Study
With the data loaded, we can assess the calibration of the foreclosure predictions using the approximate Bayesian model select approach via the bayes_ms() function. We start with assessing calibration rather than boldness-recalibration, because it is good practice to understand the level of calibration we are starting at before making any adjustments.
First, notice in the results below that the information returned in optim_details indicates successful convergence on MLE values. The MLE $\hat{\delta}_{MLE}=11.109$ indicates that the predictions need substantial shifting (on the log odds scale) to maximize their calibration. The MLE $\hat{\gamma}_{MLE}=1.271$ indicates these predictions would benefit from a slight increase in boldness to maximize calibration.
@@
While it is important to check the convergence of the MLEs to ensure reliable results, the primary goal of bayes_ms() is to assess calibration. The results above indicate that the foreclosure predictions are not well calibrated, as their posterior probability of calibration is 0.000.
We now show the likelihood ratio test for calibration implemented in llo_lrt(). As expected, the MLEs and optimization details from llo_lrt() match those from bayes_ms(). Additionally, the p-value is near zero, also indicating these predictions are not well calibrated.
@@
To quantify the maximum achievable probability of calibration of these predictions, we can first MLE recalibrate using mle_recal() and then assess the calibration of the MLE recalibrated set using bayes_ms(). This is executed in the code below and the output indicates that the maximum achievable probability of calibration of these predictions is 0.9998.
> mle_probs <- mle_recal(foreclosure\$x, foreclosure\$y, probs_only = TRUE)
> bayes_ms(mle_probs, foreclosure\$y)\$posterior_model_prob
[1] 0.9998
Now that we have a baseline value of calibration, we can explore 95%, 90%, and 80% boldness-recalibration using the brcal() function. Only the truncated output from 95% boldness-recalibration is shown below.
> br95 <- brcal(foreclosure\$x, foreclosure\$y)
@@
> br90 <- brcal(foreclosure\$x, foreclosure\$y, t=0.9, print_level=0)
> br80 <- brcal(foreclosure\$x, foreclosure\$y, t=0.8, print_level=0)
@@
Notice above that the optimization routine terminated because our stopping condition of xtol_rel = 1e-06 was met, meaning the algorithm stops when the change in parameter values from one iteration to another, relative to the size of the parameter values, is smaller than $1\times 10^{-6}$. We consider this successful convergence.
The estimated 95% boldness-recalibration parameters are $\hat{\delta}_{0.95}=12.135$ and $\hat{\gamma}_{0.95}=1.412$. Additionally, we see that the maximum achievable boldness under the constraint of $P_{\mathbf{x}^{\prime}}(M_{c}|\mathbf{y})\geq 0.95$ is $s_{b}=0.199$. Note that on average, 95% boldness-recalibration on this data takes about 33 seconds to run (on Ubuntu 22.04, R version 4.4.2).
The estimated MLEs, 95%, 90%, and 80% boldness-recalibration parameters and their achieved $P_{\mathbf{x}^{\prime}}(M_{c}|\mathbf{y})$ are visualized in Figure 4, generated by the code below. This plot supports our conclusion that both the boldness-recalibration and MLE optimizations converged appropriately, as the MLEs fall directly at the optimum of $P_{\mathbf{x}^{\prime}}(M_{c}|\mathbf{y})$ (where the graph is the deepest shade of red) and the boldness-recalibration estimates fall along the upper boundary of their respective contours.
@@
Figure 4: Color contour plot demonstrating the plot_params() function. The points in white along the contours denote the boldness-recalibration parameters. The $\times$ denotes the MLEs. The color of each grid cell represents $P_{\mathbf{x^{\prime}}}(M_{c}|\mathbf{y})$, where $\mathbf{x^{\prime}}$ is LLO-adjusted by the corresponding $\delta$ and $\gamma$ on the $x$ and $y$-axes, respectively.
Below we visualize how MLE recalibration and boldness-recalibration (for t=0.95, 0.9, and 0.8) change the individual probabilities themselves using lineplot(). The code below generates the plot shown in Figure 5. Given the size of this data, the plot may take a few minutes to generate. In practice, users can leverage the strategies discussed in Section 4.6.1 for reducing computation time.
> lineplot(foreclosure\$x, foreclosure\$y, t_levels=c(0.95, 0.9, 0.8))
Figure 5: Demonstration of the lineplot() function visualizing how predictions change from varying recalibration strategies. The first column of points is the original set of probability predictions. The second column are the MLE recalibrated predictions (i.e., after recalibrating with $\hat{\delta}_{MLE}$ and $\hat{\gamma}_{MLE}$). The last three columns are the predictions after 95%, 90%, and 80% boldness-recalibration (i.e., $\mathbf{x}_{0.95}$, $\mathbf{x}_{0.90}$, and $\mathbf{x}_{0.80}$) respectively. A line is used to connect each original prediction to where it ends up after each recalibration procedure. Blue points and lines correspond to predictions for which there was a foreclosure (i.e., “events”). Red corresponds to transactions in which there was no foreclosure (i.e., “non-events”). Achieved $P_{\mathbf{x}^{\prime}}(M_{c}|\mathbf{y})$ are reported on the x-axis in parentheses.
There are a few key takeaways from the visualization in Figure 5. First, the original foreclosure predictions are relatively accurate in terms of classification accuracy as more of the observations are colored blue (corresponding to foreclosure) for higher probabilities and more are colored red (corresponding to non-foreclosure) for lower probabilities. In fact, these predictions achieve an area under the ROC curve (AUC) of 0.8. Note that accuracy as measured by AUC does not change under any LLO-adjustment, since LLO-adjustment does not change the order of the predictions and AUC considers every possible threshold in the data. However, measures of accuracy that require a single fixed threshold to determine which observations should be classified as an “event” or “non-event” may change under LLO-adjustment, as the adjustment may push predictions to the opposite side of the threshold. Second, the original predictions are substantially biased, as they consistently underestimate the rate of foreclosure, which on average in this dataset set is 0.797. This bias is most likely due to using a model trained on pre-recession data from 2005 to estimate foreclosure rates for transactions from post-recession in 2010. MLE recalibration corrects this bias by shifting all predictions towards the observed base rate of foreclosure of 0.797. This is further reflected by the large value of $\hat{\delta}_{MLE}$, as larger values of $\delta$ indicate greater shifts in the predictions. The mean prediction shifts from 0.38 to 0.80. Finally, these predictions can also benefit from being spread out under MLE recalibration, meaning we can both increase calibration and boldness at the same time, producing predictions more useful for decision making. The spread in the predictions increases from 0.16 in the original set (range: 0.06, 0.6), to 0.18 in the MLE recalibrated set (range: 0.26, 0.95). This increase in boldness arises from the fact that $\hat{\gamma}_{MLE}>1$. Then, with a slight relaxation of calibration probability from the maximum achievable, 0.9998, to 0.95, the spread in predictions increases further to 0.20. This is the maximum boldness achievable while maintaining a calibration probability of 0.95. The further increase in spread is also reflected in the fact that $\hat{\gamma}_{0.95}>\hat{\gamma}_{MLE}>1$.
@@ 6 Conclusion
In this paper, we formulate boldness-recalibration as a nonlinear optimization. We provide details showing that the MLEs for $\delta$ and $\gamma$ maximize $P_{\mathbf{x}^{\prime}}(M_{c}|\mathbf{y})$, which provides insight for interpretation of MLE recalibration across applications. By maximizing the LLO likelihood, one maximizes $P_{\mathbf{x}^{\prime}}(M_{c}|\mathbf{y})$. We discuss various optimization routines considered for estimation of MLEs and boldness-recalibration parameters. Additionally, we present the BRcal package. The BRcal package offers a variety of tools for assessing and improving probability predictions. One key contribution of this R package is the brcal() function which allows users to maximize the boldness (spread) in predictions while maintaining a specified level of calibration as measured by the posterior model probability of calibration. The other main contributions of the BRcal package are (i) MLE recalibration, (ii) calibration assessment, and (iii) supporting visualizations. The BRcal package, with documentation and a user-friendly tutorial are available to users on CRAN and Github. Overall, the BRcal package is a useful tool to achieve more actionable predictions for better decision making. Lastly, we provide a new case study related to foreclosure predictions.
@@ Appendix A Derivation of Gradients
This appendix shows the derivation of the gradients shown in Sections A.2, A.3 and A.4 using related intermediate results presented in Section A.1.